OpenStudy (anonymous):

could somebody please check if my answer is correct. Thanks the firm would like to invest $1,000,000 with the following projected cash flow streams: Year 1: -$100,000 Year 2: -$100,000 Year 3: $100,000 Year 4: $300,000 Year 5: $1,000,000 What is the Net present value of the project at a 5% discount rate?

OpenStudy (anonymous):

PV1 = -100,000(1.05)^-1 = -95,238 PV2 = -100,000(1.05)^-2 = -90,702 PV3 = 100,000(1.05)^-3 = 86,383 PV4 = 300,000(1.05)^-4 = 246,810 PV5 = 1,000,000(1.05)^-5 = 783,526

OpenStudy (anonymous):

The NPV = 930,779

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@Hero

OpenStudy (atlas):

can u tell me why you are multiplying the net cash flow with powers of 1/1.05 instead of multipying it with powers of 0.95?

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

i used the formula PV = FV/(1+R)^n

OpenStudy (anonymous):

the discount rate is 5% = 0.05

OpenStudy (anonymous):

have a made a mistake?

OpenStudy (anonymous):

i*

OpenStudy (anonymous):

do you have another way of doing it @atlas ???

OpenStudy (atlas):

no chad you are right

OpenStudy (anonymous):

are you completely 100% sure that my answer is correct?

OpenStudy (atlas):

yes if your calculations are right

OpenStudy (anonymous):

are you familiar with these sorts of questions because i still have one to go. i hope you can help me

OpenStudy (anonymous):

do you have time?

OpenStudy (atlas):

i read these sometime back chad......i hope i can help

OpenStudy (anonymous):

give me a sec so that i can copy and paste it

OpenStudy (anonymous):

Which bond to invest in: Bond A has a Face Value of $500,000, life of 4 years, coupon rate of 6% p.a. And a yield return of 5% p.a. Bond B has a Face Value of $100,000, life of 5 years, coupon rate of 7% p.a. And a yield return of 6% p.a

OpenStudy (anonymous):

how can we find which bond is better than the other and what formula are we supposed to use?

OpenStudy (atlas):

instead of concentrating on formula - find the future value of BOND A and BOND B after 20 years (LCM of 4 and 5)

OpenStudy (anonymous):

c= FV x Coupon Rate c= 500,000 x 0.06 = 30,000

OpenStudy (anonymous):

am i right @atlas

OpenStudy (anonymous):

???

OpenStudy (atlas):

u are right but that is just the coupon rate you have to calculate the fv using yield too

OpenStudy (anonymous):

see this part i partially know

OpenStudy (anonymous):

should it be 30,000(1.05)^-1 + 30,000(1.05)^-2 + ...

OpenStudy (anonymous):

30,000(1.05)^-1 + 30,000(1.05)^-2 + 30,000(1.05)^-3 + 30,000(1.05)^-4 + 500,000(1.05)^-4 ???

OpenStudy (anonymous):

is it correct?

OpenStudy (atlas):

yeah right for BOND A

OpenStudy (atlas):

this is present value of BOND A

OpenStudy (anonymous):

now once i find the NPV OF Bond A i compare it with the NVP of Bond B? How will i know which one is better than the other?

OpenStudy (anonymous):

is it the one with the greater NPV?

OpenStudy (atlas):

no

OpenStudy (atlas):

you have to find how much return you get for per dollar invested

OpenStudy (atlas):

Calculate NPV/(invested amount) for BOND A and Bond B Whichever is greater

OpenStudy (anonymous):

can i do the first part (the one that i now understand) and then you can guide me through it

OpenStudy (atlas):

sure

OpenStudy (anonymous):

thanks @atlas

OpenStudy (anonymous):

For instance: if Bond A has a Face Value of $700,000, life of 6 years, coupon rate of 6% p.a. And a yield return of 7% p.a. it will look like this